OpenStudy (anonymous):

equation of a line perpendicular to y=2(x+5)-22, passing thru the point (-2,1)

OpenStudy (anonymous):

y - 1 = -1/2(x+2)

OpenStudy (anonymous):

what getting the y by itself? like y=???

OpenStudy (anonymous):

y = 1 + (-1/2)(x+2)

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

y = -1/2x